# Text-Tuple-Table: Towards Information Integration in Text-to-Table Generation via Global Tuple Extraction

### v1

@@ Abstract
The task of condensing large chunks of textual information into concise and structured tables has gained attention recently due to the emergence of Large Language Models (LLMs) and their potential benefit for downstream tasks, such as text summarization and text mining. Previous approaches often generate tables that directly replicate information from the text, limiting their applicability in broader contexts, as text-to-table generation in real-life scenarios necessitates information extraction, reasoning, and integration. However, there is a lack of both datasets and methodologies towards this task. In this paper, we introduce LiveSum, a new benchmark dataset created for generating summary tables of competitions based on real-time commentary texts. We evaluate the performances of state-of-the-art LLMs on this task in both fine-tuning and zero-shot settings, and additionally propose a novel pipeline called T3(Text-Tuple-Table) to improve their performances. Extensive experimental results demonstrate that LLMs still struggle with this task even after fine-tuning, while our approach can offer substantial performance gains without explicit training. Further analyses demonstrate that our method exhibits strong generalization abilities, surpassing previous approaches on several other text-to-table datasets. Our code and data can be found at https://github.com/HKUST-KnowComp/LiveSum-TTT.
@@ 1 Introduction
Reading extensive texts is demanding and time-consuming for humans, further compounded by the challenge of effectively capturing the key elements. Consequently, recent works have shifted to explore the structured summarization of text Jain et al. (2024), with tables being one highly prevalent form Wu et al. (2022); Li et al. (2023c); Sundar et al. (2024). These approaches improve text comprehension by extracting inherent yet valuable structural information from long unstructured text and enabling their applications in downstream scenarios, such as question answering Chen et al. (2020); Zhu et al. (2024), text summarization Wiseman et al. (2017); Wang et al. (2020); Mulwad et al. (2023) and text data mining Li et al. (2023b); Sui et al. (2024).
@@
Apart from the research gap at the benchmark level, in terms of methodology, considerable attention has been given to studying the ability of LLMs to comprehend and generate complex structured outputs Tang et al. (2023); Jain et al. (2024), driven by the exceptional success of LLMs in recent years Touvron et al. (2023); Anthropic (2024); OpenAI (2024). Extensive benchmarks indicate that LLMs exhibit sub-optimal performance in zero-shot settings, with multiple cases of generating inaccurate contents deviate from the given text Tang et al. (2023). To address this issue, more sophisticated prompting mechanisms have been proposed Wei et al. (2022); Khot et al. (2022); Dua et al. (2022). Among them, Jain et al. (2024) introduce a divide-and-generate prompting approach to generate more accurate and informative tables, demonstrating its effectiveness in improving model performance. However, this simplistic approach of dividing text into paragraphs and generating tables is unsuitable in more complicated situations because table-relevant information may not be contiguous in the original text and may span across various paragraphs. Therefore, developing a robust prompting method is also needed for generating complex tables that capture crucial information from scattered text or paragraphs.
To resolve the aforementioned research gaps, we introduce a novel benchmark, LiveSum, which consists of 3,771 text-based live commentaries from real-world football matches, intending to evaluate the models’ ability to generate summary tables. Unlike previous benchmarks, our benchmark necessitates the model to possess the ability to extract correct and meaningful information from complex textual data, specifically emphasizing information integration, reasoning, and conceptualization skills Wang et al. (2023a). This is because commentaries in close temporal proximity or with similar semantic meanings may describe the same event, while verbs with similar meanings may refer to the same types of events. For example, in Figure 1, the second and third dialogue boxes both describe the same goal event, and the verbs “goal” and “score” refer to the same goal event.
@@ 3.1 Live Commentary Generation
To address formatting issues in the original textual live commentary on the website, we paraphrase the text to match the commentator’s style while ensuring a certain degree of diversity. Building upon previous studies Kim et al. (2023a); Chen et al. (2023); Kim et al. (2023b), we employ ChatGPT OpenAI (2022) to generate complete live commentary automatically. Subsequently, to comply with privacy regulations and prevent bias in LLM benchmarks, we also anonymize the data leveraging named entity recognition (NER) techniques Qi et al. (2020) to produce the final version of the live commentary.
@@ 4.1 Text-to-Tuple
Considering the superior performance and flexibility of LLMs in information extraction compared to traditional techniques Ma et al. (2023); Xu et al. (2023), we employ an LLM as our tuple extractor. We follow the instructions from InstructUIE Wang et al. (2023b) and design the following prompting:
@@ 4.2 Information Integration
where <Instrction> is the directive for the current task, and <Tuples> consists of the tuples extracted in the prior stage.
@@ 4.3 Tuple-to-Table
where <Instrction> is the directive for the current task, and <Tuples> consists of the tuples produced in the prior stage.
@@ 5 Experimental Setup
In this study, we conduct fine-tuning on the LiveSum dataset using three representative open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al. (2023), LLaMA-2 Chat 7B and LLaMA-2 Chat 13B Touvron et al. (2023). We fine-tune these models following the current state-of-the-art fine-tuning methodologies Tang et al. (2023). Therefore, the outcomes represent the best results achievable with the present fine-tuning methods. We also evaluate eight state-of-the-art LLMs in zero-shot settings: LLaMA-2 Chat 13B, LLaMA-2 Chat 70B Touvron et al. (2023), Mistral Large MistralAI (2024), Claude 2.1 Anthropic (2023), Claude 3 Opus Anthropic (2024), ChatGPT OpenAI (2022), and GPT-4 OpenAI (2024). For each model, we conduct tests using two types of prompts. The first type directly describes the task by providing an instruction text $\mathbf{y}$ and accompanying it with the text $\mathbf{x}$. The second type uses the Chain-of-Thought (CoT) prompting Wei et al. (2022), incorporating the phrase “let’s think step by step” into the instruction text. See more details in Appendix C.
@@
As the generated cell content in this task consists of numerical values, we utilize commonly employed metrics in regression tasks, namely the Root Mean Square Error (RMSE). We also report the Error Rate (ER) for each cell, defining a cell as erroneous if its content does not exactly match the ground truth.
@@ 6.1 Benchmarking (RQ1)
We first analyze the performance of existing state-of-the-art LLMs on the LiveSum dataset under fine-tuning and zero-shot settings, with results displayed in Figure 4 and Table 1. Overall, the performance of most models in the zero-shot setting far exceeds that in the fine-tuning setting, indicating that the previous state-of-the-art fine-tuning method has limited capability for information integration, and there is substantial room for improvement on this benchmark. In the zero-shot setting, it is noteworthy that most models show a slight improvement in both metrics after applying CoT. Among them, the best-performing models are Mistral Large, GPT-4, and Claude 3 Opus, which are nearly comparable. They achieve RMSEs ranging from 2.08 to 2.27 and error rates between 46.20% and 48.33%. Nevertheless, this still highlights a notable deficiency in the information integration capabilities of LLMs in the zero-shot settings, underscoring the challenges and significance of our benchmark. We then analyze performance across three categories of difficulty.
@@ 6.3.2 Performance of T3 on Wiki40b Dataset
where $G(\mathcal{S})$ is the list of Question-Answer pairs $(q_{i},a_{i})$ generated by the LLM based on text $\mathcal{S}$, $Q(\mathcal{T},q)$ is the LLM’s answer to question $q$ based on table $\mathcal{T}$, and $E_{(q,a)}[x]$ is the LLM’s evaluation of whether answer $a$ and $x$ are equivalent for question $q$. On top of this evaluation, we add a step where an LLM is used to pre-screen each $(q,a)$ pair based on text $\mathcal{S}$, filtering out any pairs where the question can not be correctly answered. This process further assures the quality of the QA pairs generated by $G$. We opt for ChatGPT as the LLM for evaluation and randomly sample 500 passages for the test dataset following Jain et al. (2024). We also introduce T2(Text-Tuple) which treats the intermediary tuples from T3 as a single table $\mathcal{T}$. We aim to investigate the extent of information loss during the conversion from tuples to tables through this configuration. Figure 5 shows the Auto-QA Coverage of three methods. The curve indicates the percentage of generated tables meeting a given coverage threshold. Overall, T3 demonstrates a substantial improvement over the prior Devide-and-Generate method Jain et al. (2024). For example, when the coverage threshold is set to 70%, about 65% of data reach this threshold after applying T3, compared to only 50% with the preceding approach. It is important to note that T2 outperforms T3, with 83% of data retaining the same coverage after tuple extraction. This suggests that information loss occurs during the transformation of raw tuples into structured tables. Exploring ways to mitigate this loss represents an essential area for further research. More details are discussed in Appendix F.
@@ 7 Related Work
Many studies have been proposed to perform text-to-table generation, converting it into sequence-to-sequence problems Wu et al. (2022); Li et al. (2023c), or framing them as question-answering problems Sundar et al. (2024). With the rise of LLMs, some research has also explored evaluating LLMs under fine-tuning or zero-shot settings, and it shows that fine-tuning yields highly effective results Tang et al. (2023); Sundar et al. (2024). However, these methods employ datasets that only require the model to extract relevant information from text and populate tables, which significantly limits the scope of this task. Therefore, we introduce a new challenging dataset and propose a universal solution that greatly enhances the performance of LLMs under zero-shot setting.
@@
Information Extraction (IE) is critical and foundational for many downstream tasks in NLP. Many works have been conducted to leverage LLMs and provide effective solutions for IE tasks within a generative framework Ma et al. (2023); Lu et al. (2023); Wan et al. (2023); Zhou et al. (2024). Recent progress in LLMs also has led to the development of unified frameworks that model various IE tasks and domains Wang et al. (2023b); Sainz et al. (2024). This aligns with our intention to harness this capability to address general text-to-table generation tasks.
@@
Prompt engineering has been essential for enhancing LLMs and has demonstrated great success across a wide range of applications Wei et al. (2022); Dua et al. (2022); Li et al. (2023a); Wang et al. (2024). Among the various prompting techniques, we find the decomposed prompting Khot et al. (2022), which breaks down complex tasks into easier sub-tasks via prompting, highly effective for text-to-table generation. Jain et al. (2024) adopts this idea by breaking the text into small pieces for table generation. However, we argue that such a decomposition approach is impractical because text is not always easily divisible. For example, in our dataset, such division might result in adjacent sections describing the same event, causing errors. Hence we propose a more intuitive and broadly applicable task decomposition pipeline.
@@ Acknowledgements
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
@@ A.2 Summary Table Generation
We recruit five workers who are interested in football and are from English-speaking countries to annotate the events and employ a majority voting approach to resolve any disagreements. We randomly sample 100 cases with discrepancies and find that all errors are due to carelessness. Therefore, we conclude that the quality of the annotation quality is assured.
@@ C.1 Fine-Tune Setting
We use the open-source library named LLaMA-Factory [^2] https://github.com/hiyouga/LLaMA-Factory Zheng et al. (2024) to fine-tune all models. LoRA Hu et al. (2022) is used as the fine-tuning paradigm to accommodate our computational resources. The pre-trained weights are downloaded from the huggingface library Wolf et al. (2020). We load the models with FP16 as the precision and optimize them with an Adam optimizer Kingma and Ba (2015). The learning rate is set to 5e-5 and the batch size is 32. The maximum length for the input and generated sentence concatenation is 3,500. We warm up the model with 3,000 steps and evaluate the model every 300 steps. A linear scheduler is also used. The LoRA rank is set to 8, and the $\alpha$ is set to 32.

### v2

@@ Abstract
The task of condensing large chunks of textual information into concise and structured tables has gained attention recently due to the emergence of Large Language Models (LLMs) and their potential benefit for downstream tasks, such as text summarization and text mining. Previous approaches often generate tables that directly replicate information from the text, limiting their applicability in broader contexts, as text-to-table generation in real-life scenarios necessitates information extraction, reasoning, and integration. However, there is a lack of both datasets and methodologies towards this task. In this paper, we introduce LiveSum, a new benchmark dataset created for generating summary tables of competitions based on real-time commentary texts. We evaluate the performances of state-of-the-art LLMs on this task in both fine-tuning and zero-shot settings, and additionally propose a novel pipeline called T3(Text-Tuple-Table) to improve their performances. Extensive experimental results demonstrate that LLMs still struggle with this task even after fine-tuning, while our approach can offer substantial performance gains without explicit training. Further analyses demonstrate that our method exhibits strong generalization abilities, surpassing previous approaches on several other text-to-table datasets. Our code and data can be found at https://github.com/HKUST-KnowComp/LiveSum.
@@ 1 Introduction
Reading extensive texts is demanding and time-consuming for humans, further compounded by the challenge of effectively capturing the key elements. Consequently, recent works have shifted to explore the structured summarization of text Jain et al. (2024), with tables being one highly prevalent form Wu et al. (2022); Li et al. (2023d); Sundar et al. (2024). These approaches improve text comprehension by extracting inherent yet valuable structural information from long unstructured text and enabling their applications in downstream scenarios, such as question answering Chen et al. (2020); Zhu et al. (2024), text summarization Wiseman et al. (2017); Wang et al. (2020); Mulwad et al. (2023) and text data mining Li et al. (2023c); Sui et al. (2024).
@@
Apart from the research gap at the benchmark level, in terms of methodology, considerable attention has been given to studying the ability of LLMs to comprehend and generate complex structured outputs Tang et al. (2023); Jain et al. (2024), driven by the exceptional success of LLMs in recent years Touvron et al. (2023); Anthropic (2024); OpenAI (2024a); Jiang et al. (2023b). Extensive benchmarks indicate that LLMs exhibit sub-optimal performance in zero-shot settings, with multiple cases of generating inaccurate contents deviate from the given text Tang et al. (2023). To address this issue, more sophisticated prompting mechanisms have been proposed Wei et al. (2022); Khot et al. (2022); Dua et al. (2022). Among them, Jain et al. (2024) introduce a divide-and-generate prompting approach to generate more accurate and informative tables, demonstrating its effectiveness in improving model performance. However, this simplistic approach of dividing text into paragraphs and generating tables is unsuitable in more complicated situations because table-relevant information may not be contiguous in the original text and may span across various paragraphs. Therefore, developing a robust prompting method is also needed for generating complex tables that capture crucial information from scattered text or paragraphs.
To resolve the aforementioned research gaps, we introduce a novel benchmark, LiveSum, which consists of 3,771 text-based live commentaries from real-world football matches, intending to evaluate the models’ ability to generate summary tables. Unlike previous benchmarks, our benchmark necessitates the model to possess the ability to extract correct and meaningful information from complex textual data, specifically emphasizing information integration, reasoning, and conceptualization skills Wang et al. (2023b, a, 2024b, 2024c); Wang and Song (2024); Zheng et al. (2024a). This is because commentaries in close temporal proximity or with similar semantic meanings may describe the same event, while verbs with similar meanings may refer to the same types of events. For example, in Figure 1, the second and third dialogue boxes both describe the same goal event, and the verbs “goal” and “score” refer to the same goal event.
@@ 3.1 Live Commentary Generation
To address formatting issues in the original textual live commentary on the website, we paraphrase the text to match the commentator’s style while ensuring a certain degree of diversity. Building upon previous studies Kim et al. (2023); Chen et al. (2023); Jiayang et al. (2024); Deng et al. (2023); Jiayang et al. (2023); Chan et al. (2024b), we employ ChatGPT OpenAI (2022) to generate complete live commentary automatically. Subsequently, to comply with privacy regulations and prevent bias in LLM benchmarks Li et al. (2023a); Susnjak (2022); Li et al. (2024b); Lukas et al. (2023); Li et al. (2024a), we also anonymize the data leveraging named entity recognition (NER) techniques Qi et al. (2020) to produce the final version of the live commentary.
@@ 4.1 Text-to-Tuple
Considering the superior performance and flexibility of LLMs in information extraction compared to traditional techniques Ma et al. (2023); Xu et al. (2023), we employ an LLM as our tuple extractor. We follow the instructions from InstructUIE Wang et al. (2023c) and design the following prompting:
@@ 4.2 Information Integration
where <Instruction> is the directive for the current task, and <Tuples> consists of the tuples extracted in the prior stage.
@@ 4.3 Tuple-to-Table
where <Instruction> is the directive for the current task, and <Tuples> consists of the tuples produced in the prior stage.
@@ 5 Experimental Setup
In this study, we conduct fine-tuning on the LiveSum dataset using three representative open-source LLMs: Mistral-7B-Instruct-v0.2 Jiang et al. (2023a), LLaMA-2 Chat 7B and LLaMA-2 Chat 13B Touvron et al. (2023). We fine-tune these models following the current state-of-the-art fine-tuning methodologies Tang et al. (2023). Therefore, the outcomes represent the best results achievable with the present fine-tuning methods. We also evaluate eight state-of-the-art LLMs in zero-shot settings: LLaMA-2 Chat 13B, LLaMA-2 Chat 70B Touvron et al. (2023), Mistral Large MistralAI (2024), Claude 2.1 Anthropic (2023), Claude 3 Opus Anthropic (2024), ChatGPT OpenAI (2022), and GPT-4 OpenAI (2024a). Because of the limitation of the input lengths, we evaluate two LLMs in few-shot settings: ChatGPT-16k and GPT-4o OpenAI (2024b). For each model, we conduct tests using two types of prompts. The first type directly describes the task by providing an instruction text $\mathbf{y}$ and accompanying it with the text $\mathbf{x}$. The second type uses the Chain-of-Thought (CoT) prompting Wei et al. (2022), incorporating the phrase “let’s think step by step” into the instruction text. See more details in Appendix C.
@@
As the generated cell content in this task consists of numerical values, we utilize commonly employed metrics in regression tasks, namely the Root Mean Square Error (RMSE).
$RMSE=\sqrt{\frac{\sum_{i=1}^{n}\left(y_{i}-\tilde{y}_{i}\right)^{2}}{n}}$
where $n$ represents the number of cells, $y_{i}$ and $\tilde{y}_{i}$ represent the contents of the cell at index $i$ (for 2D tables, the index is calculated after reshaping to the 1D table) in the ground truth table and the output table, respectively.
We also report the Error Rate (ER) for each cell, defining a cell as erroneous if its content does not exactly match the ground truth.
@@ 6 Experiments and Analyses
Model Easy Medium Hard Average
RMSE/ER RMSE/ER RMSE/ER RMSE/ER
[…]
0-shot 0.22/12.0 1.76/65.1 3.22/87.2 2.08/57.3
\ \ $\Diamond$ w/ CoT 0.23/12.7 1.75/64.4 3.14/87.4 2.05/57.2
1-shot 0.25/13.5 2.29/71.9 4.33/90.4 2.76/61.9
\ \ $\Diamond$ w/ CoT 0.18/8.82 3.18/73.8 4.52/91.5 3.24/62.0
[…]
0-shot 0.00/0.13 1.51/50.4 1.82/78.5 1.43/44.8
\ \ $\Diamond$ w/ CoT 0.00/0.20 1.31/41.7 2.00/78.1 1.41/40.4
1-shot 0.00/0.19 1.24/49.0 2.19/78.8 1.44/44.3
\ \ $\Diamond$ w/ CoT 0.00/0.13 1.19/45.0 2.72/84.3 1.63/43.6
5-shots 0.00/0.20 1.71/57.2 2.15/80.0 1.65/48.6
\ \ $\Diamond$ w/ CoT 0.00/0.13 1.55/56.3 2.02/72.9 1.53/46.4
Table 2: The performance of two LLMs under the few-shot setting. We bold the best results and underline the second-best results.
@@ 6.1 Benchmarking (RQ1)
We first analyze the performance of existing state-of-the-art LLMs on the LiveSum dataset under fine-tuning, zero-shot and few-shot settings, with results displayed in Figure 4, Table 1 and Table 2. Overall, the performance of most models in the zero-shot setting far exceeds that in the fine-tuning setting, indicating that the previous state-of-the-art fine-tuning method has limited capability for information integration, and there is substantial room for improvement on this benchmark. In the zero-shot setting, it is noteworthy that most models show a slight improvement in both metrics after applying CoT. Among them, the best-performing models are Mistral Large, GPT-4, and Claude 3 Opus, which are nearly comparable. They achieve RMSEs ranging from 2.08 to 2.27 and error rates between 46.20% and 48.33%. Nevertheless, this still highlights a notable deficiency in the information integration capabilities of LLMs in the zero-shot settings, underscoring the challenges and significance of our benchmark. The results in Table 2 indicate that few-shot learning does not consistently improve model performance on this task. However, there is a significant reduction in the error rate for the hard portion under the 5-shots-CoT setting. We speculate that in such complex tasks, the model may not effectively learn and understand the examples, and instead, be influenced by the examples, leading to incorrect answers. We then analyze performance across three categories of difficulty.
@@ 6.3.2 Performance of T3 on Wiki40b Dataset
where $G(\mathcal{S})$ is the list of Question-Answer pairs $(q_{i},a_{i})$ generated by the LLM based on text $\mathcal{S}$, $Q(\mathcal{T},q)$ is the LLM’s answer to question $q$ based on table $\mathcal{T}$, and $E_{(q,a)}[x]$ is the LLM’s evaluation of whether answer $a$ and $x$ are equivalent for question $q$. On top of this evaluation, we add a step where an LLM is used to pre-screen each $(q,a)$ pair based on text $\mathcal{S}$, filtering out any pairs where the question can not be correctly answered. This process further assures the quality of the QA pairs generated by $G$. We opt for ChatGPT as the LLM for evaluation and randomly sample 500 passages for the test dataset following Jain et al. (2024). We also introduce T2(Text-Tuple) which treats the intermediary tuples from T3 as a single table $\mathcal{T}$. We aim to investigate the extent of information loss during the conversion from tuples to tables through this configuration. Figure 5 shows the Auto-QA Coverage of three methods. The curve indicates the percentage of generated tables meeting a given coverage threshold. Overall, T3 demonstrates a substantial improvement over the prior Divide-and-Generate method Jain et al. (2024). For example, when the coverage threshold is set to 70%, about 65% of data reach this threshold after applying T3, compared to only 50% with the preceding approach. It is important to note that T2 outperforms T3, with 83% of data retaining the same coverage after tuple extraction. This suggests that information loss occurs during the transformation of raw tuples into structured tables. Exploring ways to mitigate this loss represents an essential area for further research. More details are discussed in Appendix F.
@@ 7 Related Work
Many studies have been proposed to perform text-to-table generation, converting it into sequence-to-sequence problems Wu et al. (2022); Li et al. (2023d), or framing them as question-answering problems Sundar et al. (2024). With the rise of LLMs, some research has also explored evaluating LLMs under fine-tuning or zero-shot settings, and it shows that fine-tuning yields highly effective results Tang et al. (2023); Sundar et al. (2024). However, these methods employ datasets that only require the model to extract relevant information from text and populate tables, which significantly limits the scope of this task. Therefore, we introduce a new challenging dataset and propose a universal solution that greatly enhances the performance of LLMs under zero-shot setting.
@@
Information Extraction (IE) is critical and foundational for many downstream tasks in NLP. Many works have been conducted to leverage LLMs and provide effective solutions for IE tasks within a generative framework Ma et al. (2023); Lu et al. (2023); Wan et al. (2023); Zhou et al. (2024). Recent progress in LLMs also has led to the development of unified frameworks that model various IE tasks and domains Wang et al. (2023c); Sainz et al. (2024). This aligns with our intention to harness this capability to address general text-to-table generation tasks.
@@
Prompt engineering has been essential for enhancing LLMs and has demonstrated great success across a wide range of applications Wei et al. (2022); Dua et al. (2022); Li et al. (2023b); Chan et al. (2023b); Wang et al. (2024d); Yim et al. (2024); Chan et al. (2023a); Lin et al. (2024); Chan et al. (2024a). Among the various prompting techniques, we find the decomposed prompting Khot et al. (2022), which breaks down complex tasks into easier sub-tasks via prompting, highly effective for text-to-table generation. Jain et al. (2024) adopts this idea by breaking the text into small pieces for table generation. However, we argue that such a decomposition approach is impractical because text is not always easily divisible. For example, in our dataset, such division might result in adjacent sections describing the same event, causing errors. Hence we propose a more intuitive and broadly applicable task decomposition pipeline.
@@ Acknowledgements
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong.
@@ A.2 Summary Table Generation
We recruit five workers who voluntarily agreed to participate in this research without receiving any compensation to annotate the events and employ a majority voting approach to resolve any disagreements. These five workers are all postgraduate students who have . We randomly sample 100 cases with discrepancies and find that all errors are due to carelessness. Therefore, we conclude that the quality of the annotation quality is assured.
@@ C.1 Fine-Tune Setting
We use the open-source library named LLaMA-Factory [^2] https://github.com/hiyouga/LLaMA-Factory Zheng et al. (2024b) to fine-tune all models. LoRA Hu et al. (2022) is used as the fine-tuning paradigm to accommodate our computational resources. The pre-trained weights are downloaded from the huggingface library Wolf et al. (2020). We load the models with FP16 as the precision and optimize them with an Adam optimizer Kingma and Ba (2015). The learning rate is set to 5e-5 and the batch size is 32. The maximum length for the input and generated sentence concatenation is 3,500. We warm up the model with 3,000 steps and evaluate the model every 300 steps. A linear scheduler is also used. The LoRA rank is set to 8, and the $\alpha$ is set to 32.